Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below. Indicate the missing amount for each letter. Assume in all cases that manufacturing overhead is applied on the basis of direct labor cost and the rate is the same.

Case A

Case B

Case C

Direct materials used $Manufacturing cost data for Orlando Company, which (a) $91,200 $69,000
Direct labor 52,200 143,800 Manufacturing cost data for Orlando Company, which (h)
Manufacturing overhead applied 42,804 Manufacturing cost data for Orlando Company, which (d) Manufacturing cost data for Orlando Company, which (i)
Total manufacturing costs 149,800 Manufacturing cost data for Orlando Company, which (e) 216,100
Work in process 1/1/14 Manufacturing cost data for Orlando Company, which (b) 21,300 18,400
Total cost of work in process 208,600 Manufacturing cost data for Orlando Company, which (f) Manufacturing cost data for Orlando Company, which (j)
Work in process 12/31/14 Manufacturing cost data for Orlando Company, which (c) 11,900 Manufacturing cost data for Orlando Company, which (k)
Cost of goods manufactured 193,500 Manufacturing cost data for Orlando Company, which (g) 232,600
Answer #1
Concepts and reason

Goods Manufactured Schedule: The cost of goods manufactured is calculated in a statement. For the calculation of cost of goods manufactured it takes the beginning and ending work in process, direct materials, direct labour, and manufacturing overhead into consideration.

Job order Costing: It is a method of cost accounting, in which cost is collected and accumulated for each job, work order, or project separately.Especially the job order costing is followed in organizations where customized goods are produced.

Manufacturing overhead costs: The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing overhead costs or indirect costs.

Direct labour cost: the labor cost includes the salaries or wages paid to employees who help in producing goods or services. Labor refers to the actual work that the employees do to manufacture the products. Labor costs refer to the amount of money that the employer pays the employees on completion of the work.

Completion of Job: When a job is completed, the cost of the job is transferred to finished goods inventory from work-in-process inventory.

Fundamentals

The formula for direct materials used is shown below:

Directmaterialused=(TotalmanufacturingcostDirectlaborManufacturingoverheadapplied){\rm{Direct material used}} = \left( \begin{array}{l}\\{\rm{Total manufacturing cost}}\\\\ - {\rm{Direct labor}}\\\\ - {\rm{Manufacturing overhead applied}}\\\end{array} \right)

The formula for work in process in the beginning of the year (January 01, 2014) is shown below:

Workinprocess(January01,2014)=(TotalcostofworkinprocessTotalmanufacturingcost){\rm{Work in process (January 01, 2014)}} = \left( \begin{array}{c}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Total manufacturing cost}}\\\end{array} \right)

The formula for work in process at the end of the year (December 31, 2014) is shown below:

Workinprocess(December31,2014)=(TotalcostofworkinprocessCostofgoodsmanufactured){\rm{Work in process}}\left( {{\rm{December 31, 2014}}} \right) = \left( \begin{array}{l}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Cost of goods manufactured}}\\\end{array} \right)

The formula for manufacturing overhead applied is shown below:

Manufacturingoverheadapplied}=Overheadrate×Directlabor\left. \begin{array}{c}\\{\rm{Manufacturing overhead}}\\\\{\rm{ applied}}\\\end{array} \right\} = {\rm{Overhead rate}} \times {\rm{Direct labor}}

The formula for total manufacturing cost is shown below:

Totalmanufacturingcost=(Directmaterialused+Directlabor+Manufacturingoverheadapplied){\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}}\\\\ + {\rm{Direct labor}} + {\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right)

The formula for total cost of work in process is shown below:

Totalcostofworkinprocess=(Totalmanufacturingcost+Workinprocess(January01,2014)){\rm{Total cost of work in process}} = \left( \begin{array}{l}\\{\rm{Total manufacturing cost}}\\\\ + {\rm{Work in process (January 01, 2014)}}\\\end{array} \right)

The formula for cost of goods manufactured is shown below:

Costofgoodsmanufactured}=(TotalcostofworkinprocessClosingworkinprocess)\left. \begin{array}{c}\\{\rm{Cost of goods }}\\\\{\rm{manufactured}}\\\end{array} \right\} = \left( \begin{array}{l}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Closing work in process}}\\\end{array} \right)

The formula for direct labor and manufacturing overhead applied is found out together is shown below:

Totalmanufacturingcost=(Directmaterialused+Directlabor+Manufacturingoverheadapplied){\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}} + \\\\{\rm{Direct labor}} + {\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right)

Substitute manufacturing overhead applied with the below formula:

Manufacturingoverheadapplied=Overheadrate×Directlabor{\rm{Manufacturing overhead applied}} = {\rm{Overhead rate}} \times {\rm{Direct labor}}

That is,

Totalmanufacturingcost=(Directmaterialused+Directlabor+(Overheadrate×Directlabor)){\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}} + \\\\{\rm{Direct labor}} + \left( \begin{array}{l}\\{\rm{Overhead rate}} \times \\\\{\rm{Direct labor}}\\\end{array} \right)\\\end{array} \right)

The formula for manufacturing overhead applied is shown below:

Manufacturingoverheadapplied}=Overheadrate×Directlabor\left. \begin{array}{l}\\{\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right\} = {\rm{Overhead rate}} \times {\rm{Direct labor}}

The formula for total cost of work in process is shown below:

Totalcostofworkinprocess}=(Totalmanufacturingcost+Beginningworkinprocess)\left. \begin{array}{l}\\{\rm{Total cost of }}\\\\{\rm{work in process}}\\\end{array} \right\} = \left( \begin{array}{l}\\{\rm{Total manufacturing cost}}\\\\ + {\rm{Beginning work in process}}\\\end{array} \right)

(a)

Compute the direct materials used in case A:

Directmaterialused=(TotalmanufacturingcostDirectlaborManufacturingoverheadapplied)=$149,800$52,200$42,804=$54,796\begin{array}{c}\\{\rm{Direct material used}} = \left( \begin{array}{l}\\{\rm{Total manufacturing cost}}\\\\ - {\rm{Direct labor}}\\\\ - {\rm{Manufacturing overhead applied}}\\\end{array} \right)\\\\ = \$ 149,800 - \$ 52,200 - \$ 42,804\\\\ = \$ 54,796\\\end{array}

(b)

Compute the work in process (January 01, 2014) in case A:

Workinprocess(January01,2014)=(TotalcostofworkinprocessTotalmanufacturingcost)=($208,600$149,800)=$58,800\begin{array}{c}\\{\rm{Work in process (January 01, 2014)}} = \left( \begin{array}{c}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Total manufacturing cost}}\\\end{array} \right)\\\\ = \left( {{\rm{\$ 208,600}} - \$ 149,800} \right)\\\\ = \$ 58,800\\\end{array}

(c)

Compute the work in process (December 31, 2014) in case A:

Workinprocess(December31,2014)=(TotalcostofworkinprocessCostofgoodsmanufactured)=($208,600$193,500)=$15,100\begin{array}{c}\\{\rm{Work in process}}\left( {{\rm{December 31, 2014}}} \right) = \left( \begin{array}{l}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Cost of goods manufactured}}\\\end{array} \right)\\\\ = \left( {\$ 208,600 - \$ 193,500} \right)\\\\ = \$ 15,100\\\end{array}

(d)

Compute overhead rate:

It is given in case A that manufacturing overhead applied is $42,804 and direct labor is $52,200.

Overheadrate=ManufacturingoverheadappliedDirectlabor×100=42,80452,200×100=82%\begin{array}{c}\\{\rm{Overhead rate}} = \frac{{{\rm{Manufacturing overhead applied}}}}{{{\rm{Direct labor}}}} \times 100\\\\ = \frac{{42,804}}{{52,200}} \times 100\\\\ = 82\% \\\end{array}

Compute the manufacturing overhead applied in case B:

Manufacturingoverheadapplied}=Overheadrate×Directlabor=82%×$143,800=$117,916\begin{array}{c}\\\left. \begin{array}{c}\\{\rm{Manufacturing overhead}}\\\\{\rm{ applied}}\\\end{array} \right\} = {\rm{Overhead rate}} \times {\rm{Direct labor}}\\\\{\rm{ = 82\% }} \times \$ 143,800\\\\ = \$ 117,916\\\end{array}

(e)

Compute total manufacturing cost in case B:

Totalmanufacturingcost=(Directmaterialused+Directlabor+Manufacturingoverheadapplied)=($91,200+$143,800+$117,916)=$352,916\begin{array}{c}\\{\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}}\\\\ + {\rm{Direct labor}} + {\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right)\\\\ = \left( \begin{array}{l}\\\$ 91,200 + \$ 143,800\\\\ + \$ 117,916\\\end{array} \right)\\\\ = \$ 352,916\\\end{array}

(f)

Compute total cost of work in process in case B:

Totalcostofworkinprocess=(Totalmanufacturingcost+Workinprocess(January01,2014))=($352,916+$21,300)=$374,216\begin{array}{c}\\{\rm{Total cost of work in process}} = \left( \begin{array}{l}\\{\rm{Total manufacturing cost}}\\\\ + {\rm{Work in process (January 01, 2014)}}\\\end{array} \right)\\\\ = \left( {\$ 352,916 + \$ 21,300} \right)\\\\ = \$ 374,216\\\end{array}

(g)

Compute cost of goods manufactured in case B:

Costofgoodsmanufactured}=(TotalcostofworkinprocessClosingworkinprocess)=($374,216$11,900)=$362,316\begin{array}{c}\\\left. \begin{array}{c}\\{\rm{Cost of goods }}\\\\{\rm{manufactured}}\\\end{array} \right\} = \left( \begin{array}{l}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Closing work in process}}\\\end{array} \right)\\\\ = \left( {\$ 374,216 - \$ 11,900} \right)\\\\ = \$ 362,316\\\end{array}

(h)

Compute direct labor in case C:

First, calculate the manufacturing overhead cost.

Totalmanufacturingcost=(Directmaterialused+Directlabor+Manufacturingoverheadapplied){\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}} + \\\\{\rm{Direct labor}} + {\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right)

Substitute manufacturing overhead applied with the below formula:

Manufacturingoverheadapplied=Overheadrate×Directlabor{\rm{Manufacturing overhead applied}} = {\rm{Overhead rate}} \times {\rm{Direct labor}}

That is,

Totalmanufacturingcost=(Directmaterialused+Directlabor+(Overheadrate×Directlabor))$216,100=$69,000+Directlabor+82100Directlabor$216,100=$69,000+Directlabor+0.82Directlabor216,100=$69,000+1.82Directlabor\begin{array}{c}\\{\rm{Total manufacturing cost}} = \left( \begin{array}{l}\\{\rm{Direct material used}} + \\\\{\rm{Direct labor}} + \left( \begin{array}{l}\\{\rm{Overhead rate}} \times \\\\{\rm{Direct labor}}\\\end{array} \right)\\\end{array} \right)\\\\\$ 216,100 = \$ 69,000 + {\rm{Direct labor}} + \frac{{82}}{{100}}{\rm{Direct labor}}\\\\\$ 216,100 = \$ 69,000 + {\rm{Direct labor}} + 0.82{\rm{Direct labor}}\\\\216,100 = \$ 69,000 + 1.82{\rm{Direct labor}}\\\end{array}

Directlabor=$216,000$69,0001.82Directlabor=$80,824\begin{array}{c}\\{\rm{Direct labor}} = \frac{{\$ 216,000 - \$ 69,000}}{{1.82}}\\\\{\rm{Direct labor}} = \$ 80,824\\\end{array}

(i)

Compute manufacturing overhead applied in case C:

Manufacturingoverheadapplied}=Overheadrate×Directlabor=82%×$80,824=$66,276\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Manufacturing }}\\\\{\rm{overhead applied}}\\\end{array} \right\} = {\rm{Overhead rate}} \times {\rm{Direct labor}}\\\\{\rm{ = 82\% }} \times {\rm{\$ 80,824}}\\\\{\rm{ = \$ 66,276}}\\\end{array}

(j)

Compute total cost of work in process in case C:

Totalcostofworkinprocess}=(Totalmanufacturingcost+Beginningworkinprocess)=($216,100+$18,400)=$234,500\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total cost of }}\\\\{\rm{work in process}}\\\end{array} \right\} = \left( \begin{array}{l}\\{\rm{Total manufacturing cost}}\\\\ + {\rm{Beginning work in process}}\\\end{array} \right)\\\\ = \left( {\$ 216,100 + \$ 18,400} \right)\\\\ = \$ 234,500\\\end{array}

(k)

Compute the work in process (December 31, 2014) in case C:

Workinprocess(December31,2014)=(TotalcostofworkinprocessCostofgoodsmanufactured)=($234,500$232,600)=$1,900\begin{array}{c}\\{\rm{Work in process}}\left( {{\rm{December 31, 2014}}} \right) = \left( \begin{array}{l}\\{\rm{Total cost of work in process}}\\\\ - {\rm{Cost of goods manufactured}}\\\end{array} \right)\\\\ = \left( {\$ 234,500 - \$ 232,600} \right)\\\\ = \$ 1,900\\\end{array}

Ans: Part a

Direct materials used in case A is $54,796

Part b

Work in process (January 01, 2014) in case A is $58,800

Part c

Work in process (December 31, 2014) in case A is $15,100

Part d

Manufacturing overhead applied in case B is $117,916

Part e

Total manufacturing cost in case B is $352,916

Part f

Total cost of work in process in case B is $374,216

Part g

Total cost of goods manufactured in case B is $362,316

Part h

Direct labor in case C is $80,824

Part i

Manufacturing overhead cost in case C is $66,276

Part j

Total cost of work in process in case C is $234,500

Part k

Work in process (December 31, 2014) in case C is $1,900